OpenStudy (amber15):

find the simple intrest Principal $500 Rate 6.3% Time 1/2 the simple intrest is $ simplify answer write in decimal or integer

OpenStudy (anonymous):

\[S.I=\frac{ P \times R \times T }{ 100 }\] S.I=Simple Interest P=Principal R=Rate T=Time

OpenStudy (amber15):

is 15.75 correct

OpenStudy (anonymous):

yes..

OpenStudy (kenljw):

500(1+.063)^1/2=515.51 an increase of 15.51